# On Randomness in Agentic Evals

### v1

@@ 2.4.1 Quantifying Randomness in Evaluation Outcomes
“Deterministic” sampling. In theory, evals can be run deterministically with temperature 0. In practice, determinism is not achievable (Yuan et al., 2025; He and Lab, 2025): modern LLM inference engines introduce various sources of non-determinism including floating-point precision, parallelization, hardware-specific optimizations, and batching strategies. We study the extent of the problem at temperature 0 (bottom half of the table). Clearly, this variance persists even with theoretically deterministic sampling (temperature 0.0). For instance, DeepSWE-preview on nano-agent (temp zero) achieves $20.4\pm 1.0\%$ (range: 18.2%–21.4%), and Qwen3-32B on R2E-Gym achieves $22.3\pm 1.8\%$ (range: 19.8%–25.2%). Counter-intuitively, the variance never decreases and sometimes increases with temperature zero (eg 0.7% variance for Qwen3-32B x nano-agent at temperature 0.6 to 1.2% variance at temperature 0). To sum up, temperature 0 does not result in determinism.
@@ 3.2 Recommendations
To enable reliable evaluation, we recommend running multiple runs per agent under test, and estimating the performance metrics from them. The required number of runs depends on the magnitude of improvement to detect and the desired statistical power (the probability of correctly identifying a real improvement when it exists). Table 2 shows the required number of runs per agent under test for detecting different improvement magnitudes at various significance levels, assuming a normal distribution of randomness. The table presents three variance scenarios corresponding to the minimum ($\sigma=0.7\%$), median ($\sigma=1.5\%$), and maximum ($\sigma=1.8\%$) standard deviations observed across our experiments.
@@ 4.1 Randomness in Large Language Models
Recent work has identified multiple sources of non-determinism in large language models. At the infrastructure level, Yuan et al. (2025) and He and Lab (2025) demonstrate that non-associative floating point operations, rounding errors, hardware configuration, and batch size variations impact reproducibility at temperature 0. For code generation specifically, Ouyang et al. (2025) find that repeated queries yield different implementations even with greedy sampling. Prompt sensitivity represents another major source of variance. Zhuo et al. (2024); Sclar et al. (2024); Andersson et al. (2025) show that meaning-preserving changes (spacing, punctuation, example ordering, case) cause substantial performance shifts. All sources of non-determinism impact agentic evaluation scores.
@@ Appendix A Statistical Power Analysis for Determining the Number of Runs
Appendix B Inference Hyper-Parameters
@@
Appendix C Pass@k Plots

### v2

@@ Abstract
[Uncaptioned image] Inference[Uncaptioned image] Analysis[Uncaptioned image] Data
@@ 2.4.1 Quantifying Randomness in Evaluation Outcomes
“Deterministic” sampling. In theory, evals can be run deterministically with temperature 0. In practice, determinism is not achievable (Yuan et al., 2025; He & Lab, 2025): modern LLM inference engines introduce various sources of non-determinism including floating-point precision, parallelization, hardware-specific optimizations, and batching strategies. We study the extent of the problem at temperature 0 (bottom half of the table). Clearly, this variance persists even with theoretically deterministic sampling (temperature 0.0). For instance, DeepSWE-preview on nano-agent (temp zero) achieves $20.4\pm 1.0\%$ (range: 18.2%–21.4%), and Qwen3-32B on R2E-Gym achieves $22.3\pm 1.8\%$ (range: 19.8%–25.2%). Counter-intuitively, the variance never decreases and sometimes increases with temperature zero (eg 0.7% variance for Qwen3-32B x nano-agent at temperature 0.6 to 1.2% variance at temperature 0). To sum up, temperature 0 does not result in determinism.
@@ 3.2 Recommendations
To enable reliable evaluation, we recommend running multiple runs per agent under test, and estimating the performance metrics from them. The required number of runs depends on the magnitude of improvement to detect and the desired statistical power (the probability of correctly identifying a real improvement when it exists). Table 2 shows the required number of runs per agent under test for detecting different improvement magnitudes at various significance levels, assuming a normal distribution of randomness. [^1] We verified this assumption via Shapiro-Wilk tests (Shapiro & Wilk, 1965) across all 12 configuration–scaffold groups ($n=10$ runs each); 11 of 12 passed at $p\geq 0.05$. See Appendix B for details. The table presents three variance scenarios corresponding to the minimum ($\sigma=0.7\%$), median ($\sigma=1.5\%$), and maximum ($\sigma=1.8\%$) standard deviations observed across our experiments.
@@ 4.1 Randomness in Large Language Models
Recent work has identified multiple sources of non-determinism in large language models. At the infrastructure level, Yuan et al. (2025) and He & Lab (2025) demonstrate that non-associative floating point operations, rounding errors, hardware configuration, and batch size variations impact reproducibility at temperature 0. For code generation specifically, Ouyang et al. (2025) find that repeated queries yield different implementations even with greedy sampling. Prompt sensitivity represents another major source of variance. Zhuo et al. (2024); Sclar et al. (2024); Andersson et al. (2025) show that meaning-preserving changes (spacing, punctuation, example ordering, case) cause substantial performance shifts. All sources of non-determinism impact agentic evaluation scores.
@@ 5 Conclusion
Acknowledgments
This work was partially supported by the WASP program funded by Knut and Alice Wallenberg Foundation. Some computation was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS).
@@ Appendix A Statistical Power Analysis for Determining the Number of Runs
Appendix B Normality of pass@1 Scores
The power analysis in Appendix A assumes that pass@1 scores are approximately normally distributed across runs. We verified this assumption empirically using the Shapiro-Wilk test (Shapiro & Wilk, 1965), which tests the null hypothesis that the data are drawn from a normal distribution.
Table 3 reports the results for all 12 configuration–scaffold groups ($n=10$ runs each). Eleven of the 12 groups pass the test at $p\geq 0.05$. The one exception, DeepSWE-preview on nano-agent ($W=0.80$, $p=0.016$), is also the group with the smallest observed variance ($\sigma=1.0\%$). With $n=10$, the test has limited power to detect departures from normality, so passing is not a confirmation of exact normality; nonetheless, the results are consistent with the normal approximation being adequate for the purpose of power analysis.
Table 3: Shapiro-Wilk normality test results for pass@1 scores across 10 runs per configuration–scaffold group. $W$ is the test statistic; a p-value $\geq 0.05$ indicates no significant departure from normality at the 5% level.
Model Scaffold Mean Std $W$ p-value
DeepSWE-preview nano-agent 31.4% 1.0% 0.803 0.016
DeepSWE-preview r2e-gym 34.4% 1.5% 0.989 0.996
DeepSWE-preview-temp0 nano-agent 20.4% 1.0% 0.881 0.134
DeepSWE-preview-temp0 r2e-gym 19.2% 1.5% 0.923 0.383
Devstral-2 nano-agent 63.5% 1.1% 0.894 0.190
Devstral-2 r2e-gym 34.9% 1.5% 0.956 0.733
Devstral-2-temp0 nano-agent 63.8% 1.6% 0.970 0.893
Devstral-2-temp0 r2e-gym 35.4% 1.7% 0.952 0.695
Qwen3-32B nano-agent 16.4% 0.7% 0.885 0.149
Qwen3-32B r2e-gym 23.9% 1.4% 0.964 0.834
Qwen3-32B-temp0 nano-agent 16.4% 1.2% 0.962 0.809
Qwen3-32B-temp0 r2e-gym 22.3% 1.8% 0.964 0.830
Appendix C Inference Hyper-Parameters
@@
Appendix D Pass@k Plots